Hydrazine, N2H4, reacts with oxygen to form nitrogen gas and water. If 2.05 g of N2H4 reacts with excess oxygen and produces 0.450 L of N2, at 295 K and 1.00 atm, what is the percent yield of the reaction?


1
Expert's answer
2021-02-04T07:05:17-0500

% yield = (actual yield / theoretical yield)×100


actual yield by PV = nRT

1×0.450L = n×0.0821×295

n = 0.01858 moles

theoretical yield by

2.05 g N2H4 × "\frac{1moleN_2H_4}{32gN_2H_4}" ×"\frac{1moleN_2}{1moleN_2H_4}"

= 0.064 moles


% yield = (0.01858/0.064) × 100

% yield = 0.2903×100

%yield = 29.03%